How many outcomes are there in the sample space for rolling 3 dice?
There are 3 ways an even number can show up on the first die.
AND
There are 3 ways an even number can show up on the second die.
AND
There are 3 ways an even number can show up on the third die.
So since "AND" means "MULTIPLY", there are 3x3x3 or 27 ways
the three dice can all come up even. So the numerator of
the probability is 27.
We calculate the denominator of the desired probability:
There are 6 ways any number can show up on the first die
AND
There are 6 ways any number can show up on the second die
AND
There are 6 ways any number can show up on the third die
So since "AND" means "MULTIPLY", there are 6x6x6 or 216 ways
the three dice can all come up any number. So the denominator
of the probability is 216.
If you roll two fair dice what is the probability a three or six will appear?
Probability of rolling a 3 is 2/36 & the probability of rolling a 6 is 5/36; probability of rolling a 3 or 5 is 2/36 + 5/36 = 7/36.
Z Score is (x-mu)/sigma. The Z-Score allows you to go to a standard normal distribution chart and to determine probabilities or numerical values.
What is the difference between Gamma distribution and Central limit theorem?
There is abig difference between them..gamma is a distribution but central limit theorm is just like a method or technique u use to approximate gamma to another distriution which is normal....stupid
How long will it take for a car to travel 203miles when it is traveling 59mph?
Speed = Distance/Time
Rearrange the formula:
Time = Distance/Speed
Time = 203/59 = 3.440677966 hours or about 3 hours and 26 minutes.
How do you use the word probable in a sentence?
It is very probable that some of the probability questions asked on this forum were asked without any attempt to first solve them.
13/52, or 1/4.
8
What is the probability of rolling a 4 on a die?
There are 6 possible numbers that can be rolled. "4" is one of those possible numbers.
Probability is therefore 1/6.
What does rolling the dice mean in slang?
If you are rolling the dice, you are gambling, or taking a chance, or taking your chances. The suggestion is that you are letting "fate" or "luck" or "chance" decide an outcome where you could have made choices but opted not to and just "let things happen" toward some outcome.
The probability of rolling 5 40 times is 12 What is the probability of rolling 5 500 times?
Probability is a number between 0 and 1. The probability of an event cannot be 12.
What is independent events in probability concepts?
The occurrence of one event does not affect the occurrence of the other event. Take for example tossing a coin. The first toss has no affect on the outcome of the second toss, so these events are independent.
If you toss 7 coins what is the probability of getting at least 4 heads?
This is a Binomial Probability Distribution, with number of trials n=7, r = 4, 5, 6, & 7 (at least 4 is 4 or more), and probability p is 1/2 or 0.5. The related link gives the binomial probabilities and the values, which are added together. The values for r = 4, 5, 6, & 7 are 0.273, 0.164, .055, .008. These add to 0.5, so the probability of getting at least 4 heads with a coin tossed 7 times is 0.5 or 50%.
Where is the load centered if the distance from the handle to the wheel axle is 4 feet?
Assuming the load is evenly distributed from the wheel to the handle, it would be 2 feet. But if it is a wheelbarrow, there is more load over the wheel than load at the handle end when lifted, so my estimate would be 1.5 feet from the wheel.
To get the answer, divide 14 by 25, which will give you .56. You then subtract that from 1.0, and the answer is .44, or 44 %.
What are 3 areas of probable volcanic activity today in the US and why?
# Hawaii - It lies underneath a mantle "hot spot" which provides an almost constant supply of magma to the earth's surface here. # Pacific North West - The Pacific crust is being subducted underneath the North American Plate here and provides a source for magma. Part of the Pacific ring of fire. i.e. Mt St. Helen's # Alaska - Also part of the ring of fire, the Pacific plate is being subducted underneath the Laurentian (North American) plate.
What are the odds of getting three Full Houses in a row at a table of 10 playing Texas Hold'em?
If all ten players were to see showdown then the chance of one of those ten players having a full house is about 15.5%. Thus to see three full houses in a row has a chance of 15.5%^3 = 0.37%.
This may not seem like much, but with a 15.5% chance of occurring it means there is a 27% chance that 3 will happen in a row if 100 hands are played.
However, you rarely, if ever, will have all ten people play to showdown. In face you'll probably only ever have 2 or 3 players at showdown. Here we need to approximate, but let's assume for each hand played the players are playing any reasonable hand, they are particularly aggressive. This results in about an 8% chance of one of the two players getting a full house. So in 100 hands there is a 4.5% chance that a full house wins three times in a row.
What is the probability that a bottle selected at random is nonconformining?
If you are talking about mass produced bottles then the probability is very low. There would certainly be some form of quality control in the system to ensure the items fall between the tolerances required.
You would have about 1/2 of 1 % chance of living to 100 since age 100 is three standard deviations (SD) from the mean. If you are with in three SD of the mean you have 99% of the data set or probablity contained. Hence, you would be missing 1% of which half would be above age 100 and half would be below (73-27) age 46.